# Discarding Lavrentiev’s Gap in Non-automonous and Non-Convex Variational Problems

### v1

@@ 1 Introduction
    A bunch of examples to the above-mentioned theorem, as well as to our further results, are provided in Section 2.3. For a quick summary see Table 1 and Table 2.
Our most general result, namely Theorem 3, is presented in Section 2. It requires a fully anisotropic condition ($H^{\textrm{conv}}$) related to ($H^{\rm iso}_{0}$) and ($H_{\Delta_{2}}$), but unlike them, it involves a balance between $f$ and the greatest convex minorant of $f_{B}^{-}$. Since the direct verification of the conditions in Theorem 3 may be difficult in practice, we provide in Theorem 4 another condition, namely ($H^{\textrm{iso}}$), which is significantly more intuitive, easier to verify and which generalizes  ($H^{\rm iso}_{0}$). The bridge between ($H^{\textrm{iso}}$) and ($H^{\textrm{conv}}$) corresponds to Theorem 5, which also sheds new light on some conditions implying the absence of the Lavrentiev gap, already present in the literature, see [BMT, Lukas23, C-b, BCM], and partially answers a question raised in [ha-aniso].
Methods. The Lagrangians that we consider in (1.1) depend on three variables: the spacial variable $x\in\Omega$, the second variable $t$ corresponding to the values of $u(x)$, and the gradient variable $\xi$ representing the values of $\nabla u(x)$. It is known that the Lavrentiev phenomenon does not arise for autonomous integrands, namely those which do not depend on the spacial variable, see [Bousquet-Pisa]. In contrast, the $x$-dependence usually involves Lavrentiev gaps, except when a balance (or anti-jump) condition is assumed involving the spacial oscillations and the growth in the gradient variable. Such a condition is perfectly suited for all the approximations involving the convolution of $u$ with a smooth kernel ${\varrho}_{{\varepsilon}}$, see e.g. [bgs-arma, Bor-Chl, BCM, BMT, Lukas23]. More specifically, in such techniques the liminf inequality $\liminf_{{\varepsilon}\to 0}E_{\Omega}(u*{\varrho}_{\varepsilon})\geq E_{\Omega}(u)$ easily follows from the Fatou lemma. The limsup inequality $\limsup_{{\varepsilon}\to 0}E_{\Omega}(u*{\varrho}_{\varepsilon})\leq E_{\Omega}(u)$ is usually based on the Jensen inequality (here ignore the boundary condition for the moment). However, this classical approximation approach seems to fail for a Lagrangian having an essential and non-convex dependence with respect to $t$.
To overcome this difficulty, we adopt the construction presented in [Bousquet-Pisa], which relies on the classical formulation of non-parametric variational problems in terms of parametric ones. The most prominent example of such a formulation is the area functional $u\in W^{1,1}(\Omega)\mapsto\int_{\Omega}\sqrt{1+|\nabla u|^{2}}{\,{\rm d}x}$ which can be expressed as the perimeter of the subgraphs of the competing functions $u$. More generally, starting from an integrand $f$, we can associate to the corresponding energy $E_{\Omega}$ its convex modification $\widehat{E}_{\Omega\times{\mathbb{R}}}$ which is another energy that acts on the subgraphs of the functions $u$. In this new formulation, the $t$ variable from the original functional becomes an additional spacial variable. The key consequence is that the non-convex behaviour of $f$ with respect to $t$ becomes harmless and the convexity with respect to $\xi$ is enough to ensure the convexity of $\widehat{E}_{\Omega\times{\mathbb{R}}}$. The fact that the parametric formulation is a way to convexify the original one has already been exploited in the setting of $\Gamma$-convergence for $BV$-functions, see [Dal-Maso], and for the formulation of necessary conditions, see [BouFra].
To be more specific, while the original energy $E_{\Omega}$ is defined on $W^{1,1}(\Omega)$, the new energy $\widehat{E}_{\Omega\times{\mathbb{R}}}$ is defined on the set of those maps $v\in L^{\infty}(\Omega\times{\mathbb{R}})$ such that their distributional derivatives $Dv$ are finite ${\mathbb{R}}^{N+1}$-valued measures. The two energies are related by the formula
$\forall u\in W^{1,1}(\Omega)\qquad E_{\Omega}(u)=\widehat{E}_{\Omega\times{\mathbb{R}}}(1_{u})\,,$
where $1_{u}$ is the indicator function of the hypograph of $u$, see (5.1). This generalizes to any function $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{N}\to{[0,\infty)}$ the classical formulation of the non-parametric minimum area problem via the minimization of the perimeter: $\int_{\Omega}\sqrt{1+|\nabla u|^{2}}=\text{Per}(u,\Omega)$. The new energy $\widehat{E}_{\Omega\times{\mathbb{R}}}$ involves a Lagrangian $\widehat{f}:\Omega\times{\mathbb{R}}\times{\mathbb{R}}^{N+1}\to[0,\infty]$ which is homogeneous of degree $1$ and convex in the last variable, see (5.9).
$v_{n}(x,t+C_{n}|x-y|)\leq v_{n}(y,t)\,.$ (1.4)
If $v_{n}$ were the indicator function of the subgraph of a certain function $u_{n}:\Omega\to{\mathbb{R}}$, this condition would mean that the graph of $u_{n}$ lies below a family of cones with bounded apertures, which in turn is equivalent to the Lipschitz continuity of $u_{n}$. To construct such a sequence $v_{n}$, as in [Bousquet-Pisa], we introduce two scales of parameters ${\varepsilon}$ and $\delta$ and define
where $*_{x}$ refers to partial convolution just with respect to $x$, while $\alpha$ is a decreasing function depending just on $t$, see the very beginning of Section 5.3. This second term is crucial to force $t\mapsto v_{n}(x,t)$ to decrease uniformly with respect to $x$, which enables us to obtain a bound $C_{n}$ as in (1.4).
For every sequence $({\varepsilon}_{n})_{n}$ and $(\delta_{n})_{n}$ decreasing to 00, the resulting maps $v_{n}\coloneqq v_{{\varepsilon}_{n},\delta_{n}}$ converge a.e. to $v$. Moreover, the liminf inequality, namely $\liminf_{n\to\infty}\widehat{E}_{\Omega\times{\mathbb{R}}}(v_{n})\geq\widehat{E}_{\Omega\times{\mathbb{R}}}(v)$, is an easy consequence of the Reshetnyak semicontinuity theorem for functionals of measures , see Proposition 6.2. The key challenge in reaching the conclusion lies in establishing the limsup inequality of Proposition 7.1 reading
One of the main difficulties in the proof of this inequality arises from the spacial dependence of the Lagrangian $f$. Let us be more specific on this point. In Theorems 1 and 2, the anti-jump behaviour of $f$ is encapsulated either in ($H^{\rm iso}_{0}$) or ($H_{\Delta_{2}}$). Both conditions yield that under a smallness condition on $\xi$ it holds $f(x,t,\xi)\lesssim(f_{B}^{-}(t,\xi)+1)$, where $B$ is any ball in the domain (see (1.3) for the definition of $f_{B}^{-}$). In ($H^{\rm iso}_{0}$), this conclusion holds when $|\xi|\lesssim{\varepsilon}^{-\min(1,N/p)}$, where ${\varepsilon}$ is the radius of the ball $B$, while in ($H_{\Delta_{2}}$), one needs that $f_{B}^{-}(t,\xi)\lesssim{\varepsilon}^{-N}$. Our main result, i.e. Theorem 3 below, is formulated under another assumption ($H^{\textrm{conv}}$) which embraces both ($H^{\rm iso}_{0}$) and ($H_{\Delta_{2}}$). Let us stress that ($H^{\textrm{conv}}$) allows us to give a unified treatment of (1.5) not only in the different settings of Theorems 1 and 2, but also of Theorem 4 and more. See Section 2.1 for more comments on ($H^{\textrm{conv}}$). The first challenging task to establish (1.5) under the condition ($H^{\textrm{conv}}$) is an effective transfer of the anti-jump behaviour exhibited by the integrand $f$ into pertinent information for the integrand $\widehat{f}$ of the convexified energy $\widehat{E}_{\Omega\times{\mathbb{R}}}$. In particular, we present in Section 7.1 estimates for the functions $v_{{\varepsilon},\delta}$ which make it possible to exploit the anti-jump condition satisfied by $\widehat{f}$. Together with a Jensen-type inequality (Lemma 7.3), this entitles us to get a suitable bound on the measure $\widehat{f}(x,t,Dv_{{\varepsilon},\delta})$, see Lemma 7.7. To complete the proof of (1.5), we divide $\Omega$ in two regions: in the bad region, where the gradient of $u$ is large, we rely on the bound of $\widehat{f}(x,t,Dv_{{\varepsilon},\delta})$ obtained above (Proposition 7.8), while in the good one, where the gradient of $u$ is small, we apply the Reschetnyak continuity theorem (Lemma 6.1). The fact that the bad region has a small measure is crucial to conclude. Let us emphasize that in the case of an autonomous Lagrangian (i.e. without dependence on $x$) as in [Bousquet-Pisa], a much simpler Jensen-type inequality was available and this partition of $\Omega$ into a good and a bad region was unnecessary.
Once the limsup inequality (1.5) is proved, we can conclude that $\lim_{n\to\infty}\widehat{E}_{\Omega\times{\mathbb{R}}}(v_{n})=\widehat{E}_{\Omega\times{\mathbb{R}}}(v)$ (Corollary 8.1). It remains to derive from this approximating sequence for $v$ an approximating sequence for $u$. This step of the proof is similar to the corresponding one in [Bousquet-Pisa]: one can select some $s\in(0,1)$ and define functions $u_{n}$ on $\Omega$ such that $1_{u_{n}}$ coincides with the indicator function of the super level set $\{(x,t):\ v_{n}(x,t)>s\}$. The convergence of $v_{n}$ to $1_{u}$ easily implies the convergence of $u_{n}$ to $u$. That each $u_{n}$ is Lipschitz continuous is a consequence of the condition (1.4) satisfied by the maps $v_{n}$. All those arguments are provided in the proof of Proposition 8.6.
This is not the end of the story however, since the maps $u_{n}$ that we have just obtained do not necessarily agree with $\varphi$ on $\partial\Omega$. A non-trivial task is to modify $u_{n}$ to get a new map $u_{n}^{{\varphi}}$ that satisfies this boundary condition. A usual construction to achieve this goal is based on a partition of unit argument. Due to the absence of any growth condition of $f$ with respect to $\xi$ however, this approach does not work directly in our situation. Instead, we rely again on the convexified energy $\widehat{E}_{\Omega\times{\mathbb{R}}}$ to localize the problem and reduce to the case when $\Omega$ is the epigraph of a Lipschitz function. But even in that setting, the strategy followed in [Bousquet-Pisa], based on local translations, cannot be repeated because of the $x$-dependence. We rely on a different approach and modify the construction of the maps $(u_{n})_{n}$ described above by using the trick of decentered convolution, that is; the smooth kernel ${\varrho}_{\varepsilon}$ is not taken radially symmetric but decentered with respect to the origin. This entitles us to exploit the regularity of $\varphi$ and get first an approximating sequence which is uniformly close (but not necessarily equal) to the map $\varphi$ on a neighborhood of $\partial\Omega$. A final truncation argument with a cut-off function then yields an approximating sequence which does agree with $\varphi$ on the boundary.
Organization of the paper. In Section 2 we present several results on the absence of the Lavrentiev phenomenon illustrated by numerous examples. Preliminary information is presented in Section 3. Section 4 translates the consequences of the main structural and anti-jump conditions to conditions satisfied on sub-level sets of the greatest convex minorant of the infimum of $f$ over small ball, namely $(f^{-}_{B})^{**}$, see Theorem 5. Section 5 introduces the construction of the initial approximation $(v_{{\varepsilon},\delta})$. The next two sections contain the main ingredients for the convergence of convexified energies of the initial approximate sequence, i.e., ‘the liminf estimate’ for initial approximation is given in Section 6 and ‘the limsup estimate’ can be found in Section 7. In Section 8, under reduced assumptions on $f$, we establish the inner approximation (ignoring the boundary condition) in Proposition 8.6 and next we consider the boundary approximation, first for special Lipschitz domains in Section 8.2 and then for any Lipschitz sets in Section 8.3. The final proof of Theorem 3 is presented in Section 9.1 while the proofs of Theorems 24 are presented in Section 9.2.
2 Statement of the main results and their consequences
We provide a general result embracing Theorems 1 and 2. We then illustrate it by several examples that cover the classical functionals as well as new ones.
2.1 Results
The heart of our study is Theorem 3 below. It holds under a general assumption which can be seen as a balance condition between the $x$-dependence and the $\xi$-growth for a Lagrangian $f(x,t,\xi)$. This condition is formulated in terms of the greatest convex minorant of the infimum of $f$ over small balls, as in  [gszg, ha-aniso, C-b], and more recently [BCM, BMT, Lukas23]. Despite not necessarily intuitive at first, it is pretty handy and allows us to treat many situations.
@@ 2.1 Results
${\left(f^{-}_{B(x,{\varepsilon})}(t,\cdot)\right)^{**}\left(\xi\right)+|\xi|^{\max(p,N)}\leq{\mathcal{k}_{2}}{{\varepsilon}^{-N}}}\qquad\Longrightarrow\qquad f\left(x,t,\xi\right)\leq\widetilde{\mathcal{C}}_{\mathcal{k}}\left[\left(f^{-}_{B(x,{\varepsilon})}(t,\cdot)\right)^{**}\left(\xi\right)+1\right]\,,$ ($H^{\textrm{conv}}$)
where **∗ ∗ denotes the greatest convex minorant of an expression in the bracket.
@@
We are in a position to present the most general of our results. We set $p^{*}\coloneqq Np/(N-p)$ if $p<N$ and $p^{*}$ is any number larger or equal to $N$ otherwise.
@@
Given a Lagrangian satisfying  ($H^{\textrm{conv}}$), one can easily incorporate extra dependence on the second variable as described in Remark 2.3. On the other hand, if $f$ is not convex with respect to the last variable, but comparable to a convex function, one can still infer the absence of the Lavrentiev gap. More precisely – if for a Carathéodory function $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{[0,\infty)}$, there exist a function $g:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{[0,\infty)}$ satisfying the conditions of Theorem 3, a constant $C>0$, and a non-negative function $\vartheta\in L^{1}(\Omega;{\mathbb{R}})$ such that
@@
Under an isotropic or orthotropic regime, the condition ($H^{\textrm{conv}}$) is equivalent to a condition that does not involve the greatest convex minorant, namely ($H^{\textrm{iso}}$) given below. This fact is provided in Theorem 5. Consequently, one can infer the absence of the Lavrentiev phenomenon under easier to verify assumptions.
@@
If $f$ is isotropic, we assume that for every $\mathcal{k}=(\mathcal{k}_{1},\mathcal{k}_{2})\in(0,\infty)^{2}$, there exists a constant ${\mathcal{C}}_{\mathcal{k}}>0$ such that for a.e. $x\in\Omega$ and all $(t,\xi)\in(-\mathcal{k}_{1},\mathcal{k}_{1})\times{\mathbb{R}^{N}}$, for every ${\varepsilon}>0$, it holds
If $f$ admits the orthotropic decomposition, i.e., $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{[0,\infty)}$ is such that
and for every $i=1,\dots,N\,,$ the function $f_{i}:\Omega\times{\mathbb{R}}\times{[0,\infty)}\to{[0,\infty)}$ is a Carathéodory function which is convex with respect to the last variable, we suppose that each $f_{i}$ satisfies ($H^{\textrm{iso}}$).
Then, for every $u\in{\mathscr{E}_{\varphi}}(\Omega)\cap W^{1,p}(\Omega)$, there exists a sequence $(u_{n})_{n\in{\mathbb{N}}}\subset W^{1,\infty}_{\varphi}(\Omega)$ such that $u_{n}\to u$ in $W^{1,p}(\Omega)$ as $n\to\infty$ and

### v3

@@ 1 Introduction
    A bunch of examples to the theorem mentioned above, covering classical functionals as well as new ones, are provided in Section $2.$ For a quick summary see Table 1 and Table 2.
In our main result, Theorem 3, we introduce a fully anisotropic condition ($H^{\textrm{conv}}$) related to ($H^{\rm iso}_{0}$) and ($H_{\Delta_{2}}$). Still, unlike them, it involves a balance between $f$ and the greatest convex minorant of $f_{B}^{-}$ in the $\xi$-variable. More specifically, given an open ball $B$ intersecting $\Omega$ and $t\in{\mathbb{R}}$, we denote by $\left(f^{-}_{B}(t,\cdot)\right)^{**}$ the supremum of all those convex functions $g:{\mathbb{R}^{N}}\to{\mathbb{R}^{N}}$ which lie below $f^{-}_{B}(t,\cdot)$ on $\mathbb{R}^{N}$} (since $f$ and thus also $f^{-}_{B}(t,\cdot)$ are nonnegative, this supremum is well-defined). We are now in a position to define the main assumption involved in the forthcoming Theorem 3:
@@
${\left(f^{-}_{B(x,{\varepsilon})}(t,\cdot)\right)^{**}\left(\xi\right)+|\xi|^{\max(p,N)}\leq{\mathcal{k}_{2}}{{\varepsilon}^{-N}}}\qquad\Longrightarrow\qquad f\left(x,t,\xi\right)\leq\widetilde{\mathcal{C}}_{\mathcal{k}}\left[\left(f^{-}_{B(x,{\varepsilon})}(t,\cdot)\right)^{**}\left(\xi\right)+1\right]\,.$ ($H^{\textrm{conv}}$)
@@
    The condition ($H^{\textrm{conv}}$) is reminiscent of many articles in the field, compare [37, 41, 23], and more recently [14, 17, 43]. Although not necessarily intuitive at first, it is quite handy and allows us to handle many situations. Observe in particular that the length of $\xi$ is now quantified both in terms of the Euclidean norm of $\xi$ and the function $\left(f^{-}_{B(x,{\varepsilon})}(t,\cdot)\right)^{**}\left(\xi\right)$. As a consequence, we will prove that ($H^{\textrm{conv}}$) implies ($H^{\rm iso}_{0}$) (for isotropic Lagrangians) and ($H_{\Delta_{2}}$) (for $\Delta_{2}$ Lagrangians).
  2. 2.
@@
Our main result asserts that under ($H^{\textrm{conv}}$), no Lavrentiev gap may occur:
@@
Here, pp* is defined as Np/(Np)Np/(N-p) if $p<N$ while $p^{*}$ is any number larger or equal to $N$ otherwise.
@@
If $f$ is not convex with respect to the last variable, but comparable to a convex function, one can still infer the absence of the Lavrentiev gap. More precisely – if for a Carathéodory function $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{[0,\infty)}$, there exist a function $g:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{[0,\infty)}$ satisfying the conditions of Theorem 3, a constant $C>0$, and a non-negative function $\vartheta\in L^{1}(\Omega;{\mathbb{R}})$ such that
@@
It is fair to say that the direct verification of the conditions in Theorem 3 may be difficult in practice, due to the presence of the greatest convex minorant. One may wonder whether the conclusion of Theorem 3 remains true when $\left(f^{-}_{B(x,{\varepsilon})}(t,\cdot)\right)^{**}$ in ($H^{\textrm{conv}}$) is replaced by $f^{-}_{B(x,{\varepsilon})}(t,\cdot)$. Let us restrict for an instant to the isotropic situation when $f$ depends on $\xi$ only through its Euclidean norm and consider the following condition:
For every $\mathcal{k}=(\mathcal{k}_{1},\mathcal{k}_{2})\in(0,\infty)^{2}$, there exists a constant ${\mathcal{C}}_{\mathcal{k}}>0$ such that for a.e. $x\in\Omega$ and all $(t,\xi)\in(-\mathcal{k}_{1},\mathcal{k}_{1})\times{\mathbb{R}^{N}}$, for every ${\varepsilon}>0$, it holds
This condition is significantly more intuitive and easier to verify than ($H^{\textrm{conv}}$), while it substantially generalizes  ($H^{\rm iso}_{0}$) in Theorem 1. Since $f^{-}_{B(x,{\varepsilon})}(t,|\xi|)\geq\left(f^{-}_{B(x,{\varepsilon})}(t,\cdot)\right)^{**}(|\xi|)$, the assumption ($H^{\textrm{conv}}$) implies ($H^{\textrm{iso}}$). Generally speaking, however, the converse implication is not true, compare Example B.1 below. But under an isotropic or orthotropic regime, those two conditions are equivalent. This is the content of our next result.
Theorem 4.
  1. 1.
    If $f$ is isotropic and satisfies ($H^{\textrm{iso}}$), then it satisfies ($H^{\textrm{conv}}$).
  2. 2.
    If $f$ admits an orthotropic decomposition, i.e. $f:\Omega\times{\mathbb{R}}\times{\mathbb{R}^{N}}\to{[0,\infty)}$ is such that
    and for every $i=1,\dots,N\,$, the function $f_{i}:\Omega\times{\mathbb{R}}\times{[0,\infty)}\to{[0,\infty)}$ is a Carathéodory function which is convex with respect to the last variable, and satisfies ($H^{\textrm{iso}}$), then $f$ satisfies ($H^{\textrm{conv}}$).
The above statement can be seen as a counterpart of [41, Theorem 1.2] for functions having isotropic or orthotropic structure. In both cases, it is shown that a balance between $f$ and $f^{-}_{B}$ on sub-level sets of $f^{-}_{B}$ given by ($H^{\textrm{iso}}$) implies a similar balance between $f$ and $(f^{-}_{B})^{**}$ as in ($H^{\textrm{conv}}$). The main difference between Theorem 4 and [41, Theorem 1.2] is the position of the constant (we have it outside $f_{B}^{-}$ and $(f_{B}^{-})^{**}$, respectively). Our proof is based on essentially different geometrical observations. We point out that the result above also allows reformulating [17, Hypothesis (H)] or [43, Assumption 1 (a4)] in the isotropic or orthotropic setting by eliminating the ** operator.
As a simple consequence of Theorem 3 and Theorem 4, we derive the non-occurence of Lavrentiev gaps for isotropic and orthotropic functionals:
@@
We assume either that $f$ is isotropic and satisfies ($H^{\textrm{iso}}$), or that $f$ admits an orthotropic decomposition in the sense of (1.10), with each $f_{i}$ satisfying ($H^{\textrm{iso}}$). Then, for every $u\in{\mathscr{E}_{\varphi}}(\Omega)\cap W^{1,p}(\Omega)$, there exists a sequence $(u_{n})_{n\in{\mathbb{N}}}\subset W^{1,\infty}_{\varphi}(\Omega)$ such that $u_{n}\to u$ in $W^{1,p}(\Omega)$ as $n\to\infty$ and